OpenStudy (anonymous):

sin^4(theta)-cos^4(theta) =1-2cos^2(theta) how do I solve this?

hartnn (hartnn):

use this \(\Large a^2-b^2 = (a+b)(a-b) \\ a = \sin^2\theta \\ b= \cos^2\theta \)